# Membership Privacy Risks of Sharpness Aware Minimization

### v3

@@ D.2 Different Membership Inference attacks
Appendix E Datasets
@@
Appendix F Experimental setup
F.1 Models
@@ F.2 $\mathcal{I}_{ent}$ experiment
F.3 Attack setup & size of data splits
We adopt the attack setting from (Tang et al., 2022; Nasr et al., 2018) to determine the partition between training data and test data and to determine the subset of the training and test data that constitutes attacker’s prior knowledge for CIFAR-100, Purchase-100 and Texas-100 datasets. We use similar strategy to determine the data split for CIFAR-10. Specifically, the attacker’s knowledge corresponds to half of the training and test data, and the MIA success is evaluated over the remaining half.
F.4 Hyperparameter tuning and empirical validation of flatness for Sharp
@@
To verify that Sharp actually finds a sharper minima, we computed the trace of the hessian matrix using Hutchinson’s method for SGD, SAM, and Sharp models on CIFAR-100. The results are in Table˜2. Higher trace indicates a sharper minima and vice versa. The trace is the largest for Sharp and smallest for SAM.
@@ F.4.1 Ball of radius $\rho$
F.4.2 Hyperparameter tuning for CIFAR-10 & CIFAR-100
@@
F.4.3 Hyperparameter tuning for Texas-100 & Purchase-100
@@
F.4.4 Hyperparameter tuning for EyePacs
@@
Appendix G Ablation Study: Comparison of different architectures
To validate consistency across different model architectures, we report results in Table˜3 using InceptionV4  [^10] https://github.com/weiaicunzai/pytorch-cifar100/blob/master/models/ and resnet18  [^11] https://github.com/inspire-group/MIAdefenseSELENA/tree/main for CIFAR-100 and CIFAR-10. We kept our $\rho$ the same across all model architectures with value 0.1. The results are consistent with our findings that SAM tends to have higher test accuracy while having higher membership attack accuracy at the same time. Overall best attack accuracy is higher for SAM for all the cases although we find mixed findings for multi-query attack accuracy specifically.

### v4

@@ D.2 Different Membership Inference attacks
Likelihood Ratio Attack (LiRA) (Carlini et al., 2022)
LiRA is a shadow-model based single–query attack that explicitly models the distributions of a scalar score for members and non-members and then performs a likelihood ratio test. For a sample $(\mathbf{x},y)$, the attacker first defines a one-dimensional score $s(\mathbf{x},y)$ from the target model, typically the negative cross-entropy loss or the (log-)confidence on the true label $y$. Using multiple shadow models trained with and without $(\mathbf{x},y)$, the attacker estimates two score distributions: one for members (IN) and one for non-members (OUT). In practice, LiRA fits parametric Gaussians
$s(\mathbf{x},y)\mid\mathrm{IN}\sim\mathcal{N}(\mu_{\mathrm{in}},\sigma_{\mathrm{in}}^{2}),\qquad s(\mathbf{x},y)\mid\mathrm{OUT}\sim\mathcal{N}(\mu_{\mathrm{out}},\sigma_{\mathrm{out}}^{2}),$
and computes the log-likelihood ratio
$\Lambda_{\mathrm{LiRA}}(\mathbf{x},y)=\log\frac{\phi\!\left(s(\mathbf{x},y);\mu_{\mathrm{in}},\sigma_{\mathrm{in}}^{2}\right)}{\phi\!\left(s(\mathbf{x},y);\mu_{\mathrm{out}},\sigma_{\mathrm{out}}^{2}\right)},$
where $\phi(\cdot;\mu,\sigma^{2})$ denotes the Gaussian density. The LiRA decision rule is then
$I_{\mathrm{LiRA}}(\hat{p}(y|\mathbf{x}),y)=\mathds{1}\big(\Lambda_{\mathrm{LiRA}}(\mathbf{x},y)\geq\tau_{\mathrm{LiRA}}\big),$ (16)
for some threshold $\tau_{\mathrm{LiRA}}$ chosen to trade off between TPR and FPR. In the “online” variant, the attacker fits both IN and OUT distributions from shadow models; in the “offline” variant, only the OUT distribution is estimated and low likelihood under the OUT model is treated as evidence of membership.
Robust MIA (RMIA) (Zarifzadeh et al., 2024)
RMIA reframes membership inference as a calibrated hypothesis test based on a pairwise likelihood ratio between a query $(\mathbf{x},y)$ and many population samples $(\mathbf{z},y_{\mathbf{z}})$. For each pair $(\mathbf{x},\mathbf{z})$, RMIA compares how the (approximate) probability of $\mathbf{x}$ and $\mathbf{z}$ change when conditioning on the event that $\mathbf{x}$ was used to train the target model. Concretely, RMIA defines
$\mathrm{LR}(\mathbf{x},\mathbf{z})\;=\;\frac{\Pr(\mathbf{x}\mid\theta)}{\Pr(\mathbf{z}\mid\theta)}\;\Big/\;\frac{\Pr(\mathbf{x})}{\Pr(\mathbf{z})},$
where $\Pr(\cdot\mid\theta)$ denotes the target model’s likelihood and $\Pr(\cdot)$ is a population prior. Intuitively, if including $\mathbf{x}$ in the target training set fits $\mathbf{x}$ disproportionately better than many other population points $\mathbf{z}$, the ratio $\mathrm{LR}(\mathbf{x},\mathbf{z})$ becomes large. RMIA samples many $\mathbf{z}$ from the population and defines a robust membership score
$R(\mathbf{x})\;=\;\frac{1}{|n_{z}|}\sum_{\mathbf{z}}\mathds{1}\!\big(\mathrm{LR}(\mathbf{x},\mathbf{z})>\gamma\big),$
where $\gamma$ is a fixed pairwise LR threshold and $n_{z}$ is number of population (non-member) samples. The attack then declares membership if $R(\mathbf{x})$ exceeds a global threshold $\tau$; by sweeping $\tau$ one obtains a calibrated ROC curve, and for a chosen FPR one can directly pick the corresponding $\tau$. In the offline mode, all reference models are OUT models trained once on population data; in the online mode, the attacker additionally trains IN reference models that explicitly include $\mathbf{x}$ in their training set, which yields a more accurate approximation of the conditional likelihoods but is more computationally expensive.
Table 3: Attack accuracy of direct threshold MIA on SGD, Sharp, ASAM, and GSAM. Analogous to SAM, optimization methods that improve generalization (ASAM, GSAM) through finding flatter minima tend to be more prone to direct threshold attacks, while optimization that looks for sharp minima instead is more robust to MIA attack while being worse off in generalization.
Dataset Algo NN Confidence Entropy M-entropy Test Acc
CIFAR-100 SGD 76.62% 77.19% 76.61% 77.30% 80.30%
Sharp 57.62% 59.69% 57.88% 59.69% 76.14%
ASAM 78.92% 79.22% 78.86% 79.31% 81.80%
GSAM 78.63% 79.23% 79.00% 79.23% 82.16%
CIFAR-10 SGD 50.00% 59.37% 59.09% 59.51% 96.00%
Sharp 50.22% 52.86% 52.47% 52.78% 92.86%
ASAM 50.48% 61.39% 61.20% 61.32% 96.66%
GSAM 50.00% 61.46% 61.38% 61.54% 96.64%
Purchase100 SGD 66.00% 66.76% 64.78% 67.13% 85.50%
Sharp 59.58% 60.96% 58.04% 61.16% 84.31%
ASAM 66.85% 66.84% 65.39% 67.03% 85.54%
GSAM 67.45% 67.72% 66.51% 67.87% 85.82%
Texas100 SGD 59.81% 65.20% 55.74% 65.13% 50.83%
Sharp 51.11% 59.89% 53.46% 59.36% 49.97%
ASAM 60.92% 67.50% 58.80% 67.10% 53.17%
GSAM 54.89% 67.07% 57.93% 67.13% 52.04%
Appendix E Other Sharpness-aware Optimizers
In this section, we discuss other variants of SAM, namely Adaptive SAM (ASAM) (Kwon et al., 2021), Guided SAM (GSAM) (Zhuang et al., 2022), and custom designed optimizer, namely Sharp. Sharp objective is designed to explicitly find a sharper minima. The objective function of Sharp is,
This objective can be seen as minimizing the loss at current $w$ while maximizing the loss in the vicinity. We empirically verify that this objective does lead to a sharper minima measuring its hessian trace. Results and discussion about Sharp are available in Section˜H.4.
The results on CIFAR10, CIFAR100, Purchase100, and Texas100 are reported in Table 3. Other sharpness-aware optimizers are shown to achieve similar generalization gain, albeit at the cost of higher membership attack accuracy. On the other hand, optimizer that explicitly looks for a sharp minima does worse in terms of generalization, but has better membership privacy.
Table 4: Comparison of offline shadow model MIA on SGD and SAM. In blue we highlight the best performing model on the test set, and in red the model with higher privacy leakage (higher AUC, Attack Accuracy, and TPR@0.1%FPR).
SGD SAM
Dataset Attack Test Acc AUC Attack Acc TPR@.1 Test Acc AUC Attack Acc TPR@.1
CIFAR-100 RMIA 67.7% 86.8% 77.3% 17.3% 69.1% 87.7% 78.1% 18.9%
LiRA 76.2% 71.9% 16.9% 77.8% 73.2% 19.4%
CIFAR-10 RMIA 92.3% 69.4% 62.3% 4.3% 93.1% 72.7% 64.6% 5.7%
LiRA 54.1% 55.9% 4.1% 58.7% 58.3% 7.0%
Purchase100 RMIA 76.5% 68.7% 62.6% 1.7% 77.4% 70.2% 63.6% 1.9%
LiRA 52.9% 53.4% 0.1% 53.7% 54.2% 0.1%
Texas100 RMIA 46.9% 74.8% 67.4% 3.6% 49.2% 78.8% 69.8% 5.6%
LiRA 56.9% 58.3% 0.8% 61.7% 61.5% 2.6%
Appendix F Offline Shadow Model Attacks
We report the results for offline shadow model attacks in Table Table˜4. The results are in line and support the finding that SAM tends to incur higher membership privacy leakage. Excluding cases for tabular datasets where TPR at 0.1% FPR is near zero for both models, SAM has higher values for all other attack metrics. Consistent with the literature, RMIA is more effective for offline setting compared to LiRA (Zarifzadeh et al., 2024). For online setting, we use a different experimental setup (WideResNet targets and shadows, our own training pipeline, and a different choice of auxiliary z-points). In this setting, LiRA is slightly stronger than RMIA across most metrics (Table 3), but the gaps are modest and much smaller than those reported for the offline comparison in Zarifzadeh et al. (2024). We therefore view our results as broadly consistent with prior work: RMIA and LiRA are competitive state-of-the-art shadow-model attacks, and their exact ranking can depend on architectural and training choices. Our main conclusions—in particular, that SAM consistently increases vulnerability to both LiRA and RMIA compared to SGD—are unaffected by these small differences.
Appendix G Datasets
@@
Appendix H Experimental setup
H.1 Models
@@ H.2 $\mathcal{I}_{ent}$ experiment
H.3 Attack setup & size of data splits
We adopt the attack setting from (Tang et al., 2022; Nasr et al., 2018) to determine the partition between training data and test data and to determine the subset of the training and test data that constitutes attacker’s prior knowledge for CIFAR-100, Purchase-100 and Texas-100 datasets. We use similar strategy to determine the data split for CIFAR-10. Specifically, the attacker’s knowledge corresponds to half of the training and test data, and the MIA success is evaluated over the remaining half. For shadow model attacks, the total sample pool used is 50000 for CIFAR10 and CIFAR100, 40000 for Purchase100, and 20000 for Texas100. For RMIA attack, we used $\gamma=1$ and selected all of the $z$ samples within the training pool that were not part of the target model’s training set. On CIFAR10, for example, number of $z$ samples was 25000.
H.4 Hyperparameter tuning and empirical validation of flatness for Sharp
@@
To verify that Sharp actually finds a sharper minima, we computed the trace of the hessian matrix using Hutchinson’s method for SGD, SAM, and Sharp models on CIFAR-100. The results are in Table˜5. Higher trace indicates a sharper minima and vice versa. The trace is the largest for Sharp and smallest for SAM.
@@ H.4.1 Ball of radius $\rho$
H.4.2 Hyperparameter tuning for CIFAR-10 & CIFAR-100
@@
H.4.3 Hyperparameter tuning for Texas-100 & Purchase-100
@@
H.4.4 Hyperparameter tuning for EyePacs
@@
Appendix I Ablation Study: Comparison of different architectures
To validate consistency across different model architectures, we report direct threshold attack results in Table˜6 using InceptionV4  [^11] https://github.com/weiaicunzai/pytorch-cifar100/blob/master/models/ and resnet18  [^12] https://github.com/inspire-group/MIAdefenseSELENA/tree/main for CIFAR-100 and CIFAR-10. We kept our $\rho$ the same across all model architectures with value 0.1. The results are consistent with our findings that SAM tends to have higher test accuracy while having higher membership attack accuracy at the same time. Overall best attack accuracy is higher for SAM for all the cases although we find mixed findings for multi-query attack accuracy specifically.
@@ Appendix I Ablation Study: Comparison of different architectures
C10: RMIA (Online)
C10: RMIA (Offline)
C10: LiRA (Online)
C10: LiRA (Offline)
C100: RMIA (Online)
C100: RMIA (Offline)
C100: LiRA (Online)
C100: LiRA (Offline)
Purch: RMIA (Online)
Purch: RMIA (Offline)
Purch: LiRA (Online)
Purch: LiRA (Offline)
Texas: RMIA (Online)
Texas: RMIA (Offline)
Texas: LiRA (Online)
Texas: LiRA (Offline)
Figure 7: ROC curves comparing SAM (Orange) vs. SGD (Blue) across all datasets and attack modes on log-log scale. Rows represent datasets (CIFAR-10, CIFAR-100, Purchase100, Texas100). Columns represent the attack configuration. The ROC curve for SAM (orange) is above the ROC curve for SGD (blue) for nearly the entire range for most settings.